OpenStudy (curry):

Find all harmonic polynomials u(x,y) of degree three, that is, u(x,y) = ax^3 + bx^2y + cxy^2 + dy^3

OpenStudy (dumbcow):

sorry im not familiar with these, I think the Laplacian needs to be zero which has to do the gradient but i havn't worked with these im guessing this is diff equ or advanced calc 3

OpenStudy (dumbcow):

@ganeshie8 , @jim_thompson5910 , @SolomonZelman maybe they know more about harmonic polynomial

OpenStudy (solomonzelman):

nope, I am unfamiliar with harmonic polynomials, if I had a little more time, I would google it, but... sorry-:(

OpenStudy (amistre64):

how is a harmonic poly defined? or is that what is given already?

OpenStudy (amistre64):

the google on this doesnt ring any bells for me

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Harmonic_polynomial http://en.wikipedia.org/wiki/Laplacian#Two_dimensions ok from this i think we got something take 2nd partial derivatives for x and y , then it equal to 0

OpenStudy (curry):

that's all the question gives.

OpenStudy (dumbcow):

\[\frac{du^2}{dx} = 6ax +2by\] \[\frac{du^2}{dy} = 6dy +2cx\] \[\Delta f = 6dy+2cx + 6ax +2by = 0\] \[(c+3a)x + (b+3d)y = 0\]

OpenStudy (curry):

And the chapter doesn't give much about harmonic stuff.

OpenStudy (curry):

ye, i have that much so far too. But is that what defines all harmonic polynomials?

OpenStudy (dumbcow):

i dont know im just trying to figure it out like you haha

OpenStudy (curry):

yes, i ran into that one as well

OpenStudy (dumbcow):

so we can conclude c = -3a b = -3d sub that into original and theres your answer for general form

OpenStudy (curry):

so would that be the answer for all harmonic polynomials also?

OpenStudy (dumbcow):

well all of degree 3

OpenStudy (curry):

hmm, so we need to find it for degree 2,1, and 0 also?